What speed would a 2,000-π‘˜π‘” sport utility vehicle have to attain in order to have 𝑖) the same momentum? 𝑖𝑖) the same kinetic energy?


1
Expert's answer
2021-12-10T12:30:23-0500

Suppose that the momentum is "p" and the kinetic energy is "E" . So,


"p=mv\to v=p/m=p/2000" .


"E=mv^2/2\to v=\sqrt{2E/m}=\sqrt{2E/2000}=\sqrt{E/1000}" .